# Decoding Gray Matter: large-scale analysis of brain cell morphometry to inform microstructural modeling of diffusion MR signals

### v1

@@ 1 Introduction
Here we aim to correct this imbalance with a comprehensive analysis of GM cellular morphology, looking at structural and topological morphology, and shape descriptors of over 3000 real three dimensional reconstructions from mouse, rat, monkey, and human brain cortex.
@@ 2.1 Microscopy Dataset
In order to measure characteristic features for specific cell types, the open access repository Neuromorpho.org [34] was used, for comprehensive range of cellular reconstructions available. We downloaded and analysed three dimensional reconstructions of 3,448 brain cells, from mouse, rat, monkey, and human, in the form of SWC files. The ”SWC” encodes for the last names of its initial designers Ed Stockley, Howard Wheal, and Robert Cannon and is an ASCII text-based file that describes three-dimensional neuronal or glial morphology. The SWC file defines a set of labelled nodes connected by edges characterizing the three-dimensional structure of each cell Fig.1.
The cells acquired were representative of eight cell types: microglia, astrocyte, pyramidal, granule, purkinje, glutamatergic, gabaergic, and basket cells, from mouse/rat (N=1,278), monkey (N=948), and human (N=1,222), all healthy controls.
@@ 3.2 Comparing neuronal and glial cells
Given the growing interest in differentiating neurons and glial cells, the mean values of the structural features were computed across only neuronal and glial cells and reported in Tab.3. The finding show compared to, glial cells, neurons had larger soma and cell domain and longer branches; and reduced branching, branch curvedness, number of primary projections, and proportion of soma surface covered by projections; the remaining features displayed similar values.
@@ 3.4 Distribution of some structural features of interest for dMRI modelling
The full distributions of values for some structural features of interest, obtained by merging together all the estimated values from all the species for each cell type, are shown in Fig.6.
Figure 6: Probability density functions of soma radius, branch length, intra-branch residence time and intra-cellular residence time for each cell types and all the species. The residence times were calculated from the surface to volume ratios with a permeability of $2*10^{-2}\mu m/ms$ [57] and diffusivity typical of water.
@@ 4.1 A Quantitative View of Gray Matter Microstructure
Neural Cell Membrane Permeability to Water. Several works suggest water exchange between unmyelinated neurites and ECS and/or soma occurring on time scales comparable to typical dMRI clinical acquisitions, i.e.  10-100 ms. Although there is not a consensus yet, some works report exchange times tex for ex vivo mouse brain  5-10 ms [58, 59, 30], suggesting membrane permeability  125 $\mu$m/s (like red blood cells), others report tex for in vivo mouse brain of  20-40 ms [57, 29, 60], suggesting membrane permeability  2-35 $\mu$m/s. Using the latter estimates of membrane permeability, and our estimates of the cellular and branch surface-to-volume ratios we estimated distribution of intra-domain and intra-branch residence times that span from a few milliseconds to tens of milliseconds (Fig.6). On average, glial cells have shorter residence times than neurons, and small neurons like granule cells have much shorter residence times than large neurons like pyramidal cells. These observations support the idea of a wide distribution of exchange times within the GM, with water within glial cells in faster exchange with ECS than that within neurons (10-20 ms versus 20-50 ms). We have to acknowledge here that in our analysis we did not include dendritic spines, boutons and glial leflets, that can occupy up to $20\%$ of the GM volume [61]. Recent studies have demonstrated that diffusion-mediated exchange between cellular projections’ shaft and these small lateral protrusions is another mechanism of water exchange, different from the permeative one, but happening on similar time scales (from a few millisecond to a few tens of milliseconds) [62, 63, 64]. Therefore, the interplay between diffusion-mediated and permeative exchange cannot be neglected when analysing and intepreting time-dependent dMRI measurements in the GM. Appropriate modelling and acquisition strategies to disentangle the two mechanisms and provide unbiased estimates are still missing, and represent an exciting avenue for future research in the field.
@@ 4.2 General Implications for Biophysical Modelling
  • Impact of projection curvedness is generally negligible for both water and metabolites: the impact of curvedness can only be significant when (from [66] $2D\Delta,2D\delta\geq(<R_{c}>_{s})^{2}$, that is for $\Delta,\delta\geq 750\ ms$ for water and $\Delta,\delta\geq 3750\ ms$ for metabolites, given $<R_{c}>_{s}\approx 55\ \mu m$, which is far longer than the values typically. This is further supported by [41].
@@ 4.3 Lessons Learnt for Gray Matter Microstructure Imaging through dMRI
In general, exchange cannot be completely ignored, nor does exchange fully average out the restriction effects of cell membranes, supporting including both neurite exchange and a soma compartment in a parsimonious biophysical model of diffusion in GM, as proposed by Olesen et al. [30] with the SANDIX model. There are however some caveats to bear in mind when interpreting the model estimates: (a) the soma apparent MR radius is unavoidably an overestimation of the true soma radius; (b) the compartmental signal fractions reflect, apart from volume, also unknown T2 and T1 weighting, which might differ amongst the compartments; (c) neurite’s fast exchange could be problematic for models based on the Kärger’s model because it could violate the model’s assumptions (i.e. barrier-limited diffusion: $t_{ex}\gg d^{2}/D$, where $d$ is the neurite diameter; and $\delta\ll t_{ex}$) and lead to biased estimates; (d) using single diffusion encoding acquisitions it is impossible to disentangle restriction and exchange; more refined acquisitions could allow the separation and more accurate quantification of these two effects [59, 79]. The estimates of the model parameters should therefore only be taken as an indication of the true tissue features and validation against realistic numerical simulations and/or alternative measurements in controlled phantoms (e.g. biomimetic tissues and brain organoids) and/or post-mortem samples (e.g., optical, confocal or electron microscopy) remains essential.

### v2

@@ 1 Introduction
Here we aim to correct this imbalance with a comprehensive analysis of GM cellular morphology, looking at structural and topological morphology, and shape descriptors of over 3,500 real three-dimensional reconstructions from mouse, rat, monkey, and human brain cortex.
@@ 2.1 Microscopy Dataset
In order to measure characteristic features for specific cell types, the open access repository Neuromorpho.org [34] was used, for comprehensive range of cellular reconstructions available. We downloaded and analysed three dimensional reconstructions of 3,598 brain cells, from mouse, rat, monkey, and human, in the form of SWC files. The ”SWC” encodes for the last names of its initial designers Ed Stockley, Howard Wheal, and Robert Cannon and is an ASCII text-based file that describes three-dimensional neuronal or glial morphology. The SWC file defines a set of labelled nodes connected by edges characterizing the three-dimensional structure of each cell Fig.1.
The cells acquired were representative of eight cell types: microglia, astrocyte, pyramidal, granule, purkinje, glutamatergic, gabaergic, and basket cells, from mouse/rat (N=1,358), monkey (N=948), and human (N=1,292). Although the NeuroMorpho database contains over 270,000 cellular reconstructions, only 4,278 satisfied our inclusion criteria: healthy controls; having complete reconstruction of all the dendrites; containing full information about diameters and angles; and being three-dimensional reconstructions. Of these, 680 did not pass our visual inspection on the quality of the reconstruction and were excluded. Our quality assessment criteria include: consistent estimates of dendritic diameters (e.g., most of the 680 reconstructions had a fixed nominal diameter for all the branches instead of the real one); continuity of the cellular processes; minimal reconstruction artifacts (e.g., most of the 680 reconstructions had artifactual shrinkage in the Z direction).
@@ 3.2 Comparing neuronal and glial cells
Given the growing interest in differentiating neurons and glial cells, the mean values of the structural features were computed across only neuronal and glial cells and reported in Tab.3. The findings show, compared to glial cells, neurons had larger soma and cell domain and longer branches; and reduced branching, branch curvedness, number of primary projections, and proportion of soma surface covered by projections; the remaining features displayed similar values.
@@ 3.4 Distribution of some structural features of interest for dMRI modelling
The full distributions of values for some structural features of interest, obtained by merging together all the estimated values from all the species for each cell type, are shown in Fig.6. Given the increasing interest of the microstructure imaging through dMRI community in estimating glia microstructure and the striking difference between the neuronal components of cerebral and cerebellar cortices (e.g., Purkinje cells only in cerebellum), we decided to group them into three classes: glia, cortical neurons and cerebellar neurons, and use three different colors to simplify the visualization of the results.
Figure 6: Probability density functions of soma radius, branch length, intra-branch residence time and intra-cellular residence time for each cell types and all the species. The residence times were calculated from the surface to volume ratios with a permeability of 20 $\mu m/s$ [57] and diffusivity typical of water.
@@ 4.1 A Quantitative View of Gray Matter Microstructure
Neural Cell Membrane Permeability to Water. Several works suggest water exchange between unmyelinated neurites and ECS and/or soma occurring on time scales comparable to typical dMRI clinical acquisitions, i.e.  10-100 ms. Although there is not a consensus yet, some works report exchange times $t_{ex}$ for ex vivo mouse brain  5-10 ms [58, 59, 30], suggesting membrane permeability  125 $\mu$m/s (like red blood cells), others report $t_{ex}$ for in vivo mouse brain of  20-40 ms [57, 29, 60], suggesting membrane permeability  2-35 $\mu$m/s. Using the latter estimates of membrane permeability, and our estimates of the cellular and branch surface-to-volume ratios we estimated distribution of intra-domain and intra-branch residence times that span from a few milliseconds to tens of milliseconds (Fig.6). On average, glial cells have shorter residence times than neurons, and small neurons like granule cells have much shorter residence times than large neurons like pyramidal cells. These observations support the idea of a wide distribution of exchange times within the GM, with water within glial cells in faster exchange with ECS than that within neurons (10-20 ms versus 20-50 ms). We have to acknowledge here that in our analysis we did not include dendritic spines, boutons and glial leflets, that can occupy up to $20\%$ of the GM volume [61]. Recent studies have demonstrated that diffusion-mediated exchange between cellular projections’ shaft and these small lateral protrusions is another mechanism of water exchange, different from the permeative one, but happening on similar time scales (from a few millisecond to a few tens of milliseconds) [62, 63, 64]. Therefore, the interplay between diffusion-mediated and permeative exchange cannot be neglected when analysing and intepreting time-dependent dMRI measurements in the GM. Appropriate modelling and acquisition strategies to disentangle the two mechanisms and provide unbiased estimates are still missing, and represent an exciting avenue for future research in the field.
@@ 4.2 General Implications for Biophysical Modelling
  • Impact of projections curvedness is generally negligible for both water and metabolites: the impact of curvedness can only be significant when (from [66] $2D\Delta,2D\delta\geq(<R_{c}>_{s})^{2}$, that is for $\Delta,\delta\geq 750\ ms$ for water and $\Delta,\delta\geq 3750\ ms$ for metabolites, given $<R_{c}>_{s}\approx 55\ \mu m$, which is far longer than the values typically. This is further supported by [41].
@@
  • Impact of projections orientation dispersion can be significant. Cells with highly oriented and polarized projections, such as Purkinje cells, granule cells and pyramidal neurons have high FA ($>0.50$) and low orientation dispersion ($<0.25$); while the projections of most of the glial cells and other neuronal cells have low FA ($<0.50$) and high orientation dispersion ($>0.25$). This supports growing evidence that FA, e.g., from DTI, and orientation dispersion estimates, e.g. from NODDI, can discriminate different cytoarchitectural domains (or layers) in hippocampus, cortex and cerebellum [68, 69, 70, 71, 72].
@@ 4.3 Lessons Learnt for Gray Matter Microstructure Imaging through dMRI
In general, exchange cannot be completely ignored, nor does exchange fully average out the restriction effects of cell membranes, supporting including both neurite exchange and a soma compartment in a parsimonious biophysical model of diffusion in GM, as proposed by Olesen et al. [30] with the SANDIX model. There are however some caveats to bear in mind when interpreting the model estimates: (a) the soma apparent MR radius is unavoidably an overestimation of the true soma radius; (b) the compartmental signal fractions reflect, apart from volume, also unknown T2 and T1 weighting, which might differ among the compartments; (c) neurite’s fast exchange could be problematic for models based on the Kärger’s model because it could violate the model’s assumptions (i.e. barrier-limited diffusion: $t_{ex}\gg d^{2}/D$, where $d$ is the neurite diameter; and $\delta\ll t_{ex}$) and lead to biased estimates; (d) molecular diffusion in structural features such as dendritic spines and glial leaflets can lead to diffusion-mediated exchange mechanisms that occur on the same time scales of permeative exchange, making the interpretation of exchange measurements through MRI solely due to permeative exchange fundamentally wrong; (e) using single diffusion encoding acquisitions it is impossible to disentangle restriction and exchange; more refined acquisitions could allow the separation and more accurate quantification of these two effects [59, 84]. The estimates of the model parameters should therefore only be taken as an indication of the true tissue features and validation against realistic numerical simulations and/or alternative measurements in controlled phantoms (e.g. biomimetic tissues and brain organoids) and/or post-mortem samples (e.g., optical, confocal or electron microscopy) remains essential.
4.4 Limitations
Our investigation, while thorough, is not exhaustive (nor could it be). Other features could certainly be measured and tabulated. However, we focus here on those deemed relevant for biophysical modeling in the current literature on microstructure imaging via dMR. We will release our code openly and freely to allow future work to complement this study with additional features and information as needed. Our focus is on a selected set of real three-dimensional reconstructions, as we aim to characterize cellular morphology in the healthy brain. Future studies can use the released code to incorporate more and improved reconstructions of healthy brain tissue as they become available (e.g., through updates to NeuroMorpho or large electron microscopy studies [85, 86]). We provide a few illustrative examples to demonstrate how this study can inform biophysical modeling of dMR signals. However, the information obtained from our investigation has broader applications, and we hope the scientific community will find it valuable for advancing our understanding of gray matter microstructure as a whole. Throughout the manuscript, we discussed the limitations of the morphometric approach used (e.g., an expected underestimation of $<20\%$ for soma volume and branch radius). Within the constraints of currently available tools, we provide reference values that were previously unavailable. We have taken great care to report error estimates and uncertainties for all measurements to account for these limitations. Future work is needed to improve morphometric algorithms, but this is beyond the scope of the current study.